### v2

# Powerformer: A Section-adaptive Transformer for Power Flow Adjustment

Abstract
In this paper, we present a novel transformer architecture tailored for learning robust power system state representations, which strives to optimize power dispatch for the power flow adjustment across different transmission sections. Specifically, our proposed approach, named Powerformer, develops a dedicated section-adaptive attention mechanism, separating itself from the self-attention used in conventional transformers. This mechanism effectively integrates power system states with transmission section information, which facilitates the development of robust state representations. Furthermore, by considering the graph topology of power system and the electrical attributes of bus nodes, we introduce two customized strategies to further enhance the expressiveness: graph neural network propagation and multi-factor attention mechanism. Extensive evaluations are conducted on three power system scenarios, including the IEEE 118-bus system, a realistic 300-bus system in China, and a large-scale European system with 9241 buses, where Powerformer demonstrates its superior performance over several baseline methods.
keywords:
graph neural network, power flow adjustment, power system state representation learning, section-adaptive attention, transformer
journal: xxx
Chen et al.: Powerformer: A Section-adaptive Transformer for Power Flow Adjustment
Graph Abstract:
[…]
1 Introduction
In real-world scenarios, power system operators typically focus on monitoring the operational state of a power system that involves numerous variables, placing emphasis on transmission sections rather than isolated components. As shown in Fig. 1, a transmission section consists of a group of transmission lines with the same color. To ensure power system security margins, operators commonly employ the total transfer capability of transmission sections to monitor power flows and regulate the operational state de2024discrete ; avraam2023operational ; jiang2022steady . Specifically, overloading a critical transmission section poses a serious threat to the system, potentially triggering a chain reaction of failures that may cause severe cascading blackouts. Hence, the efficient management of power flow at transmission sections is paramount, serving as a fundamental pillar to ensure the stability, resilience, and reliability of the power system.
Figure 1: The illustration of 10 transmission sections $\{\Phi_{1},...,\Phi_{10}\}$ in the IEEE 118-case system, where each section is represented by a set of transmitting lines that share the same color and are closely located electrically. An illustrative section $\Phi_{7}$ is the set of purple lines, i.e., $\Phi_{7}=\,${bus33-bus37, bus19-bus34, bus30-bus38, bus23-bus24}, that partitions the system into distinct left and right components.
To address power flow adjustment across different transmission sections, power dispatch is an effective strategy that enables precise finetuning of generation outputs to optimize power flow distribution; traditional dispatch methods, which rely on the system model, can be primarily categorized into two distinct classes: optimization-based capitanescu2011state ; qiang2023enhancing and sensitivity-based dutta2008optimal ; lai2021vulnerability techniques. Optimization-based methods utilize numerical techniques to transform the dispatch problem into a constrained programming problem, while sensitivity-based methods apply iterative calculations to find candidate generators by calculating sensitivity indices. However, these model-based methods often encounter the computational burdens and struggle to effectively control unpredictable system coupling within a limited time. Thus, traditional methods are unsuitable for effectively managing real-time control in a complex and large-scale power system.
Recently, deep reinforcement learning (DRL) has emerged as an innovative alternative to traditional power dispatch methods. This data-driven approach effectively utilizes the robust capabilities of the pertained deep neural networks, providing a powerful solution for various power flow adjustment. However, among the existing DRL methods li2023deep ; tang2023energy ; deng2024novel ; liu2023ask , there are two evident limitations in the state representations: 1) the absence of efficient awareness regarding transmission sections, which hinders the optimal performance for distinct section adjustment; 2) the high coupled electrical factors and the ignored topology structure of power system, ignificantly affect the extraction of the state characteristics.
In this paper, we strive to develop an advanced transformer architecture for learning robust power system state representations, facilitating precise power dispatch strategies that stabilize the power system under various section settings. To tackle the first limitation, we develop a section-adaptive attention mechanism to enable an efficient amalgamation of section information with the power system state, leveraging the comprehensive capabilities of the transformer architecture vaswani2017attention ; zhang2023knowledge ; zhang2022bootstrapping for the specific sectional system analysis. For the second limitation, we begin by analyzing the graph topology of the power system, and employing graph neural networks xu2018powerful ; Li2023MessagepassingST ; VERDONE2024122151 for feature propagation to embed structural information into state representations. Furthermore, we introduce a multi-factor attention mechanism to efficiently fuse each state factor, ensuring comprehensive and accurate extraction of system states. Finally, we utilize the proposed Powerformer, integrating the training of Dueling DQN Wang2015DuelingNA to achieve precise and efficient control of power flow across transmission sections, thereby leading to improvements in the overall system performance.
Our main contributions can be summarized as follows:
  • 1.
    We develop a section-adaptive attention mechanism that effectively integrates power system state features and section power flow information to achieve the knowledge amalgamation for different transmission sections.
  • 2.
    We introduce graph neural networks propagation alongside a multi-factor attention mechanism, specifically tailored to fuse the power system’s graph topology information and effectively emerge electrical state factors.
  • 3.
    The extensive evaluation on three distinct power system cases, including the IEEE 118-bus system, a realistic 300-bus system in China, and a large-scale 9241-bus European system, demonstrates our method’s superiority over several baselines with significantly improved performance.
2 PRELIMINARY
2.1 Problem Formulation and Task Setting
2.1.1 Problem Formulation
Transmission section power flow adjustment involves the regulation and control of electrical power flow through the specific section or interface of a power transmission network. In a specific power system, there exist $N$ transmission sections $\{\Phi_{1},\Phi_{2},...,\Phi_{N}\}$, where each section $\Phi_{i}$ consists of several transmission lines that not only share the same direction of active power flow but also maintain close electrical proximity to one another liu2023transmission . To accurately determine the total power flow of a transmission section $\Phi_{i}$, the combined power contributions are aggregated and meticulously calculated:
$\mathcal{P}_{\Phi_{i}}=\sum_{j\in\Phi_{i}}\mathcal{P}_{j}\quad\,{s.t.}\;\mathcal{P}_{\text{min},\Phi_{i}}\leq\,\mathcal{P}_{\Phi_{i}}\,\leq\mathcal{P}_{\text{max},\Phi_{i}},$ (1)
$\mathcal{Q}_{\Phi_{i}}=\sum_{j\in\Phi_{i}}\mathcal{Q}_{j}\quad{s.t.}\;\mathcal{Q}_{\text{min},\Phi_{i}}\leq\mathcal{Q}_{\Phi_{i}}\leq\mathcal{Q}_{\text{max},\Phi_{i}},$
where $\mathcal{P}_{\Phi_{i}}$ and $\mathcal{Q}_{\Phi_{i}}$ represent the total active and reactive power flow in the $i$-th transmission section, and $\mathcal{P}_{j}$ and $\mathcal{Q}_{j}$ denote the active and reactive power flow in the $j$-th transmission line within the section $\Phi_{i}$. $\mathcal{P}_{\text{min},\Phi_{i}}$, $\mathcal{P}_{\text{max},\Phi_{i}}$, $\mathcal{Q}_{\text{min},\Phi_{i}}$, $\mathcal{Q}_{\text{max},\Phi_{i}}$ are the minimum and maximum allowable active and reactive power flow in the $i$-th transmission section. These limits are essential for ensuring the power system operates safely and efficiently, maintaining power flow within each transmission section at levels that prevent overloading and stability issues. To address this problem, power dispatch is a commonly employed strategy for regulating power system flow and maintaining safety margin constraints.
2.1.2 Task Setting
In the context of power flow adjustment across various transmission sections, multi-task learning zhang2018overview ; TAN2023121177 is suitable to manage power dispatch for simultaneously addresses different scenarios within the power system, where each section’s power flow treated as a distinct task managed by a unique adjustment strategy. In this study, we establish the 10-section power flow adjustment task, selected based on the modularity of communities to reflect the importance of the sections, in order to evaluate the model’s performance on three distinct power system cases.
2.2 Sectional Power System Network
The power system topology, being non-Euclidean chen2020covariance ; Chen2022RiemannianLM , is typically represented as a graph structure chen2023hybrid ; sun2022deep ; chen2023tele ; chen2023improving . Therefore, we organize power system data in the form of graph-structured data. In a power system graph network, nodes represent buses and edges symbolize transmission lines. Specifically, a power system network with section information can be defined as:
Definition 1 (Sectional Power System Network) A sectional power system network is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E},\boldsymbol{H},\boldsymbol{Z}_{\Phi})$, where $\mathcal{V}=\{v_{1},...,v_{n}\}$ and =$\mathcal{E}=\{e_{1},...,e_{m}\}$ represent the sets of $n$ nodes and $m$ edges, respectively. $\boldsymbol{H}=[\boldsymbol{h}_{1},...,\boldsymbol{h}_{n}]\in\mathbb{R}^{d\times n}$ denotes the state feature matrix, where $\boldsymbol{h}_{i}$ denotes the $i$-th node. $\boldsymbol{Z}_{\Phi}$ is the representation of transmission section $\Phi$.
To illustrate a sectioned power system network, Fig. 1 provides an example with 10 sections in the IEEE 118-case system. In this case, transmission section $\Phi_{7}$ is represented by the purple lines, namely $\Phi_{7}=${bus33-bus37, bus19-bus34, bus30-bus38, bus23-bus24}. This specific section divides the system into left and right components.
2.3 Transformer Architecture
Numerous Pre-trained Language Models (PLMs) employ a multi-layer Transformer architecture vaswani2017attention for the purpose of encoding text representations. In each layer of the Transformer, a multi-head self-attention mechanism is utilized to generate contextualized representations for individual text tokens. Let $\boldsymbol{H}^{(l)}=[\boldsymbol{h}_{1}^{(l)},...,\boldsymbol{h}_{n}^{(l)}]$ denote the output feature of the $l$-th Transformer layer, where $\boldsymbol{h}_{i}^{(l)}\in\mathbb{R}^{d}$ is the hidden representation of the text token at position $i$. Subsequently, in the $(l+1)$-th Transformer layer, the multi-head self-attention (MHA) is calculated as follows:
[…]
$\text{head}^{t}(\boldsymbol{H}^{(l)})=\boldsymbol{V}_{t}^{(l)}\cdot\text{softmax}\left(\frac{{\boldsymbol{K}_{t}^{(l)\top}\boldsymbol{Q}_{t}^{(l)}}}{\sqrt{d}}\right),$ (3)
$\boldsymbol{Q}_{t}^{(l)}=\boldsymbol{W}_{Q,t}^{(l)}\boldsymbol{H}^{(l)},\;\boldsymbol{K}_{t}^{(l)}=\boldsymbol{W}_{K,t}^{(l)}\boldsymbol{H}^{(l)},\;\boldsymbol{V}_{t}^{(l)}=\boldsymbol{W}_{V,t}^{(l)}\boldsymbol{H}^{(l)},$ (4)
[…]
3 Method
3.1 Overall Architecture
An overview of the Powerformer architecture is presented in Fig. 2b, illustrating the fundamental components crucial for obtaining robust representations in sectional power system networks, such as state factorization, Graph Neural Network (GNN) propagation, and section-adaptive attention, etc. Different from the conventional Transformer, our architecture utilizes the set $\{\boldsymbol{H},\boldsymbol{Z}_{\Phi},\mathcal{E}\}$ as input, which consists of the state feature matrix, section power flow, and graph structure information. The section-specific information $\boldsymbol{Z}_{\Phi}$ is utilized to learn Query representation, enabling the development of section-adaptive attention mechanism. To address the specific electrical factors on buses, we disentangle the system state representation into $k$ distinct factors to prevent feature coupling during propagation, and then incorporate the section-adaptive attention matrices to establish the multi-factor section-adaptive attention mechanism. Furthermore, to enhance the expressiveness of Key, Value propagation, we utilize the popular GNN architecture to embed the graph structure information $\mathcal{E}$ into the final state representation.
Figure 2: The illustration of the Transformer and Powerformer architectures. (a) Transformer architecture with the multi-head self-attention mechanism (MHA). (b) Powerformer architecture with the multi-factor section-adaptive attention mechanism (MFSA).
3.2 Powerformer for Sectional Power System
State Factorization and GNN propagation. Each node in the power system includes important state factors, such as active power, reactive power, voltage magnitude, and phase angle, providing comprehensive insights into the state characteristic of power system.
If the state matrix is directly inputted into the propagation layer, it will result in the coupling of different factors, causing the redundancy in the learned features and ultimately diminishing the expressive power. Thus, we begin by disentangling the input state feature into $k=4$ factors, with the $l$-th layer’s representation of the $t$-th factor denoted by $\boldsymbol{H}_{t}^{(l)}\in\mathbb{R}^{d\times n}$.
Furthermore, for the propagation of the Key and Value components in relation to the $t$-th factor, we utilize the widely-used GNN to incorporate the power system’s graph topological structure for efficient updates:
$\boldsymbol{K}_{t}^{(l)}=\mathcal{F}_{K,t}^{(l)}(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),\;\boldsymbol{V}_{t}^{(l)}=\mathcal{F}_{V,t}^{(l)}(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),$ (5)
where $\mathcal{F}_{K,t}^{(l)},\mathcal{F}_{V,t}^{(l)}$ represent the GNN propagation to update the Key and Value: $\boldsymbol{K}_{t}^{(l)},\boldsymbol{V}_{t}^{(l)}\in\mathbb{R}^{d\times n}$. Specifically, this propagation is implemented using a Graph Isomorphism Network (GIN) model xu2018powerful :
$\boldsymbol{h}_{v}^{(l)}={\rm MLP}^{(l)}((1+\epsilon^{(l)})\cdot\boldsymbol{h}_{v}^{(l-1)}+\sum\nolimits_{u\in\mathcal{N}(v)}\boldsymbol{h}_{u}^{(l-1)}),$ (6)
where $\boldsymbol{h}_{v}^{(l)}$ represents the feature vector of node $v$ at the $l$-th layer. $\mathcal{N}(v)$ refers to the neighborhood set of node $v$ based on the edge set $\mathcal{E}$. MLP is the Multi-layer Perceptron network.
Section-adaptive Attention (SA). For section-adaptive representations, we first utilize the power flow information of the transmission lines on the current section to encode the section, and set the power flow information that is not within the section to zero. The resulting initialization representation is $\boldsymbol{Z}_{\Phi}$. Then, we use a MLP network to learn the Query component and align its dimensions with $\boldsymbol{K}_{t}^{(l)}$ and $\boldsymbol{V}_{t}^{(l)}$:
$\boldsymbol{Q}^{(l)}=\mathcal{M}_{Q}^{(l)}(\boldsymbol{Z}_{\Phi}^{(l)}),$ (7)
where $\mathcal{M}_{Q}^{(l)}$ is the employed MLP to update section representation. The updated Query is $\boldsymbol{Q}^{(l)}\in\mathbb{R}^{d}$. Moreover, it is worth noting that the Query component in our architecture does not require the state factorization. Then, the section-adaptive attention without softmax function within the $t$-factor can be expressed as:
$\boldsymbol{a}_{t}^{(l)}=\text{SA}^{t}(\boldsymbol{K}_{t}^{(l)},\boldsymbol{Q}^{(l)})=\frac{{\boldsymbol{K}_{t}^{(l)\top}\boldsymbol{Q}^{(l)}}}{\sqrt{d}},$ (8)
where the achieved $\boldsymbol{a}_{t}^{(l)}\in\mathbb{R}^{n}$ is the section-adaptive attention matrix of the $t$-th factor at the $l$-layer.
Multi-factor Section-adaptive Attention (MFSA). In order to achieve an effective representation fusion of multiple factors, it is essential to effectively integrate the weights allocated to different factors across nodes. Particularly, we recombine a new attention matrix, which is actually formed by combining different section-adaptive attention matrices, and then apply the softmax function on the dimensions of these factors:
$\widetilde{\boldsymbol{A}^{(l)}}=\text{softmax}(\boldsymbol{A}^{(l)}),$ (9)
$\boldsymbol{A}^{(l)}=[\boldsymbol{a}_{1}^{(l)},...,\boldsymbol{a}_{t}^{(l)},...,\boldsymbol{a}_{k}^{(l)}],\;\,t=1,...,k,$ (10)
where $\widetilde{\boldsymbol{A}^{(l)}}\in\mathbb{R}^{n\times k}$ is the resulting matrix by applying the softmax function along the factor’s dimension. Finally, the Multi-factor Section-adaptive Attention (MFSA) mechanism:
[…]
where $\boldsymbol{V}_{t}^{(l)}$ and $\widetilde{\boldsymbol{A}^{(l)}}$ are obtained from Eq. 5 and 9, $\widetilde{\boldsymbol{A}^{(l)}}_{t}$ denotes $t$-th column associated with $t$-th electrical factor. Symbol $\odot$ denotes the Hadamard product with broadcast mechanism. It is important to note that the left matrix side of the $\odot$ operation is in $\mathbb{R}^{d\times n}$, while the right matrix is in $\mathbb{R}^{n}$. The resulting matrix after the Hadamard product will also reside in the space $\mathbb{R}^{d\times n}$.
3.3 The Implementation of Powerformer
At the implementation stage, we merge the Eq. 8, 9 and 10 into Eq. 11 in order to provide a more clear formulation of our proposed Powerformer architecture:
[…]
$\text{factor}^{t}(\boldsymbol{H}^{(l)},\boldsymbol{Z}_{\Phi}^{(l)},\mathcal{E})=\boldsymbol{V}_{t}^{(l)}\odot\text{softmax}\left(\boldsymbol{A}^{(l)}\right)_{t},$ (13)
$\boldsymbol{A}^{(l)}=[\boldsymbol{a}_{1}^{(l)},...,\boldsymbol{a}_{t}^{(l)},...,\boldsymbol{a}_{k}^{(l)}],\;\,\boldsymbol{a}_{t}^{(l)}=\frac{{\boldsymbol{K}_{t}^{(l)\top}\boldsymbol{Q}^{(l)}}}{\sqrt{d}},$ (14)
$\begin{split}&\boldsymbol{Q}^{(l)}=\mathcal{M}_{Q}^{(l)}(\boldsymbol{Z}_{\Phi}^{(l)}),\\ \boldsymbol{K}_{t}^{(l)}=\mathcal{F}_{K,t}^{(l)}&(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),\;\boldsymbol{V}_{t}^{(l)}=\mathcal{F}_{V,t}^{(l)}(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),\end{split}$ (15)
where $\mathcal{M}_{Q}$, $\mathcal{F}_{K,t}^{(l)}$, $\mathcal{F}_{V,t}^{(l)}$ are the corresponding propagation function for the Query, Key and Value components. Notably, our section-adaptive attention based representation (Eq. 13) diverges from the self-attention based representation (Eq. 3) by employing softmax on the attention matrix (Eq. 14) that considers all factors, rather than the attention matrix of a single head. Furthermore, we employ summation to amalgamate diverse factor representations, rather than concatenation.
Complexity Analysis. The time complexity of Powerformer can be analyzed based on Multi-factor Section-adaptive Attention (MFSA) and Graph Neural Network (GNN) propagation. MFSA is $O(nd)$, where $n$ represents the number of bus nodes and $d$ represents the dimensionality of the embeddings. On the other hand, GNN propagation is $O(nd^{2}+md)$, where $m$ represents the number of transmission lines. As a result, Powerformer exhibits a time complexity of O($nd+nd^{2}+md$).
Joint Training. In order to facilitate the ongoing training of the proposed Powerformer framework, we perform iterative updates on the state representation using a deep reinforcement learning framework. Specifically, in this study, we employ the Dueling DQN method Wang2015DuelingNA to achieve this purpose.
In summary, our Powerformer approach enhances the traditional Transformer’s expressiveness for sectional power system power flow adjustment. This is achieved by incorporating the section information into attention matrix, and utilizing GNN and multi-factor mechanism to enable more effectively system state representations.
4 Case Study
To demonstrate the effectiveness of our proposed Powerformer, comprehensive case studies have been conducted. In this section, we first provide the details of experimental setting. Subsequently, we conduct a comparative analysis of our method with concat and attention-based representation learning methods, as well as several popular adjustment baselines. Finally, we conduct ablation studies to enhance the understanding of our proposed method.
4.1 Experimental Setting
Scenario Generation. To obtain adequate scenarios for training and testing purposes, we first employ two small-scale systems, namely the IEEE 118-bus system and a realistic China 300-bus system. Then, we adopt a very large European 9241-bus system from the PEGASE project into our operations Josz2016ACPF . The detailed scenario generation is summarized as follows:
  • 1.
    To simulate a variety of scenarios within the power system more effectively, we implement random perturbations to loads and generators. For each scenario, we randomly select 25% of generators and loads and then introduce random disturbances to their initial power flow ranging from 10% to 200%, with $10\%$ intervals.
  • 2.
    The power range for each transmission section $\Phi$ is pre-scheduled, as detailed in Table 1. Insecure scenarios, where the power flow of the corresponding transmission sections exceeds pre-scheduled power range, are considered for case studies during each power flow adjustment.
Table 1: The pre-scheduled lower and upper power bound (MW) of each transmission section in the IEEE 118-bus system, the realistic China 300-bus system, and the European 9241-bus system.
Section $\Phi$ 118-bus System 300-bus System 9241-bus System
Lower Upper Lower Upper Lower Upper
1 90 640 140 1000 15 110
2 50 360 280 1960 840 5880
3 40 290 170 1200 145 1000
4 90 640 240 1680 55 390
5 70 480 460 3200 560 3920
6 45 300 200 1400 360 2520
7 130 880 200 1400 345 2410
8 55 390 480 3360 760 5320
9 130 880 170 1200 130 910
10 90 615 80 590 180 1260
Finally, a total of 1829 scenarios were generated for the IEEE 118-bus system, with 1656 scenarios used for training and 173 for testing. Similarly, the China 300-bus system resulted in 1817 scenarios, including 1637 for training and 180 for testing. The European 9241-bus system produced 2037 scenarios, with 1848 for training and 189 for testing.
Figure 3: The visualization highlights the distribution of Concat, Attention, and Powerformer operations for combining state and transmission section representations on the IEEE 118-bus system.
Refer to caption (a) Concat Refer to caption (b) Attention Refer to caption (c) Powerformer (Ours)
Table 2: The performance of our method and baselines on three different power systems is evaluated based on an average of 5 trials. The performance differences between our method and baselines are indicated in parentheses, with better performance denoted by a higher test success rate and lower test economic cost. Method 118-bus System 300-bus System 9241-bus System Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Concat 68.44 (-28.17) 626,215  (+4,017) 57.66 (-37.14) 992,048 (+22,366) 37.57  (-5.29) 324,485        (-2) Attention 85.23 (-11.38) 625,128  (+2,930) 85.30  (-9.50) 972,573  (+2,891) 31.13 (-11.73) 325,152    (+665) Powerformer (Ours) 98.19 622,198 97.36 969,682 68.25 324,487
Figure 3: The visualization highlights the distribution of Concat, Attention, and Powerformer operations for combining state and transmission section representations on the IEEE 118-bus system.
Compared Methods and Parameter Settings. We first consider two basic architectures for state representations learning on sectional power system: 1) Concat architecture. We employ a MLP network that takes the concatenated vector of the task encoding vector and state feature vector as input. 2) Attention architecture. We utilize the soft attention network Yang2020MultiTaskRL , which allows agent to dynamically select different network modules based on the section representation. Furthermore, the proposed Powerformer is compared with several popular DRL methods, including Deep Q Network (DQN) MnihKSRVBGRFOPB15 , Double DQN Hasselt2015DeepRL , Dueling DQN Wang2015DuelingNA , Advantage Actor Critic (A2C) Mnih2016AsynchronousMF , Proximal Policy Optimization (PPO) schulman2017proximal .
The proposed Powerformer employs a two-layer networks configuration, specifically the hidden dimension $d$ is 64. Additionally, the section encoder utilizes a two-layer MLP network with a candidate hidden dimension set {128, 64}. To implemente the Powerformer, we utilize the Before Transformer architecture min2022transformer , a type of graph transformer. The Dueling architecture includes a two-layer value network with a dimension of 128 and a one-layer advantage network with a dimension of 128. The section representation is $\boldsymbol{Z}_{\Phi}\in\mathbb{R}^{4m}$, where 4 is the corresponding electrical factors of active power, reactive power, voltage magnitude, and phase angle, $m$ denotes the number of the transmission lines. All network modules in the architecture employ the ReLU activation function. During each training iteration, batches of 64 episodes are sampled from a replay buffer containing 20,000 episodes. The target update interval is set to 100, and the discount factor is 0.9. For exploration, we apply an epsilon-greedy approach, where epsilon linearly decreases from 0.1 to 0.01 over 500,000 training steps and remains constant thereafter.
Evaluations. The evaluation of policies incorporates metrics such as the mean test success rate, mean test economic cost, and mean inference speed across a range of tasks.
4.2 Comparison with Different State Representations
In this subsection, our primary focus revolves around a comprehensive comparison between our novel Powerformer and two commonly used representation learning methods, i.e., concat and attention-based representations. By conducting a comprehensive analysis using the same reinforcement learning framework of Dueling DQN Wang2015DuelingNA , we aim to clearly showcase the remarkable effectiveness and superiority of our proposed Powerformer method.
Figure 4: The learning curves of all methods for 4-section task and 10-section task on three power systems. The experimental results use the median performance and one standard deviation (shaded region) over 5 random seeds to ensure a fair comparison.
Refer to caption (a) Refer to caption (a) IEEE 118-bus System(4-section Task) Refer to caption (b) Realistic 300-bus System(4-section Task) Refer to caption (c) European 9241-bus System(4-section Task) Refer to caption (d) IEEE 118-bus System(10-section Task) Refer to caption (e) Realistic 300-bus System(10-section Task) Refer to caption (f) European 9241-bus System(10-section Task)
Table 3: The performance of our method and baselines in 4-section and 10-section tasks on three different power systems is evaluated based on an average of 5 trials. The performance differences between our method and baselines are indicated in parentheses, with better performance denoted by a higher test success rate and lower test economic cost. Method 118-bus System (4-section Task) 300-bus System (4-section Task) 9241-bus System (4-section Task) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) DQN 88.97 (-19.66) 693,272  (+1,130) 79.05 (-20.68) 981,695 (+19,605) 52.04 (-16.93) 325,331     (-99) Double DQN 87.63 (-12.04) 693,653  (+1,511) 80.15 (-19.58) 995,038 (+32,948) 55.41 (-13.56) 325,464     (+34) Dueling DQN 91.17  (-8.50) 692,378    (+236) 71.45 (-28.28) 995,279 (+33,189) 52.48 (-16.49) 325,266    (-164) A2C 84.98 (-14.69) 697,853  (+5,711) 93.71  (-6.02) 974,455 (+12,365) 53.07 (-15.90) 325,093    (-337) PPO 26.67 (-73.00) 673,399 (-18,743) 46.88 (-52.85) 937,454 (-24,636) 39.18 (-29.79) 326,614  (+1,184) OPF 83.33 (-16.34) 639,782 (-52,360) 56.78 (-42.59) 956,120  (-5,970) 64.47  (-4.50) 324,995    (-435) Powerformer (Ours) 99.67 692,142 99.73 962,090 68.97 325,430 Method 118-bus System (10-section Task) 300-bus System (10-section Task) 9241-bus System (10-section Task) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) DQN 61.23 (-35.38) 626,808  (+4,610) 56.74 (-38.06) 991,762 (+22,080) 13.67 (-29.19) 324,658    (+164) Double DQN 53.28 (-43.33) 638,079 (+15,881) 61.24 (-33.49) 990,126 (+20,444) 16.75 (-26.11) 324,792    (+305) Dueling DQN 68.44 (-28.17) 626,215  (+4,017) 57.66 (-37.14) 992,048 (+22,366) 37.57  (-5.29) 324,485        (-2) A2C 80.14 (-16.47) 626,703  (+4,505) 75.61 (-19.19) 976,467  (+6,785) 12.70 (-30.16) 325,000    (+513) PPO 24.28 (-72.33) 596,315 (-25,883) 48.61 (-46.19) 930,581 (-39,101) 19.40 (-23.46) 325,549 (+1,062) OPF 72.25 (-24.36) 606,259 (-15,939) 50.56 (-44.24) 953,349 (-16,333) 51.32  (+8.46) 321,586 (-2,901) Powerformer (Ours) 98.19 622,198 97.36 969,682 68.25 324,487
Figure 4: The learning curves of all methods for 4-section task and 10-section task on three power systems. The experimental results use the median performance and one standard deviation (shaded region) over 5 random seeds to ensure a fair comparison.
The performance results presented in Table 2 showcases a comparison between the Powerformer and two fundamental state encoding methods—Concat and Attention—across three different power system cases. The adopt evaluation is based on two metrics: success rate and economic cost, averaged over five trials. Among these results, it is worth noting that the proposed Powerformer demonstrates the exceptional success rates of 98.19%, 97.36%, and 68.25% respectively, solidifying its position as the standout performer. This is particularly evident in its remarkable performance when applied to large-scale power systems such as the European 9241-bus system. In terms of economic cost, Powerformer tends to be more efficient or on par with the other methods. This suggests that Powerformer can achieve higher performance without a relatively low test economic cost, an important consideration for practical applications.
Furthermore, we utilize the t-SNE toolkit to visually illustrate the distribution of state representations for 10 distinct section tasks on the IEEE 118-bus system in Fig. 3. Our proposed Powerformer approach can effectively cluster 10 different section tasks into 10 distinct central points, while the other two methods cannot accomplish. This also explains the poor performance of the other two state encoding methods, as they couple the transmission section representation and different electrical factors of node features together (Fig. 3a and 3b), making it impossible to execute specific actions.
4.3 Comparison with Different Adjustment Baselines
To demonstrate the superiority of our method in efficiently and effectively adjusting transmission section power flow, we thoroughly analyze and compare our method with various popular approaches in deep reinforcement learning (DRL) framework, as well as a well-established traditional optimal power flow (OPF) adjustment method. Moreover, in order to thoroughly validate the effectiveness of our Powerformer, we design the additional 4-section power flow adjustment tasks, which aims to deal with the top four ranked sections in terms of transmission capacity among a total of ten sections.
Through analyzing the learning curve in Fig. 4, our proposed method has successfully achieved the best test success rate among all considered architectures on three different systems. Although DQN and its variants require fewer timesteps to achieve reasonable results, Dueling DQN stands out with the highest achieved test success. This also provides advantageous conditions for our method to obtain the best results. We can analyze and compare the superiority of our method based on the difficulty levels of different tasks, especially in the more challenging task of 10-section adjustment (i.e., Fig. 4d4e and 4f), where the effectiveness of our method is highlighted. Furthermore, from the perspective of different system cases, our method performs even better on large-scale datasets. This is because our method is able to learn better representations of sectional power system states, thus avoiding the coupling of state features. We can comparatively analyze the learning curve on the IEEE 118-bus and European 9241-bus systems, and the significant differences in the effectiveness of our method can be directly observed because when the system case is simple, the task is also relatively simple, hence the superiority is not so apparent.
From a quantifiable perspective of comparison adjustment algorithms, we leverage the comparative results of success rates and economic costs pertaining to 4-section and 10-section adjustment tasks, as evidenced in Table 3, to serve as prime examples for our comprehensive analysis. Specifically, we further analyze the results by examining different power system cases. For the IEEE 118-bus system, Powerformer outperforms all other methods with a success rate of 99.67%, significantly surpassing Dueling DQN (91.17%) on the 4-section task. As complexity increases in the 10-section task, Powerformer maintains a success rate of 98.19%, outperforming A2C (80.14%) by a substantial margin. For the China 300-bus system, Powerformer continues to exhibit a satisfying level of performance in comparison to the baseline approaches, accomplishing an impressively high test success rate of $99.73\%$ and $97.36\%$ on the 4-section and 10-section tasks, respectively. For the European-9241 system, our method consistently demonstrates exceptional performance, achieving an impressive success rate of 68.97% and 68.25% when considering the test success rates of 4-section and 10-section tasks. In contrast, alternative reinforcement learning methods fail to surpass the 35% threshold for the complex 10-section task. Additionally, our method consistently demonstrates its exceptional ability to achieve relatively low economic costs across all scenarios.
Figure 5: The inference time comparison of total 173 scenarios on the IEEE 118-bus system.
Table 4: The inference time of our method and OPF method on three different power systems. 118-bus System 300-bus System 9241-bus System OPF 44.822 $\pm$ 38.459 47.668 $\pm$ 28.230 1912.608 $\pm$ 1636.746 Powerformer (Ours) 0.078 $\pm$ 0.151 0.121 $\pm$ 0.176 0.585 $\pm$ 0.602
Figure 5: The inference time comparison of total 173 scenarios on the IEEE 118-bus system.
4.4 Inference Time Analysis
In real-time applications, such as the European 9241-bus system, inference time plays a critical role, particularly in large-scale systems. In such scenarios, power flow adjustment becomes more complex due to the numerous variations that need careful consideration and strategic decision-making. To better understand the superiority of our method in terms of inference speed, we provide Fig. 5 and Table 4 to demonstrate the advantages of our approach when compared to traditional OPF method used in practical scenarios. As shown in Fig. 5, the experimental results demonstrate that, when compared to the traditional OPF method, the time required to successfully achieve the predefined range for each individual test is significantly reduced across a comprehensive set of 173 test scenarios in the IEEE 118-bus system. Moreover, it is notable that our method exhibits a noteworthy reduction in inference time across three different power systems, as evidenced by the compelling numerical statistics presented in Table 4. As a result, our proposed method is remarkably more time efficient than the traditional OPF approach, especially as the system size increases, making the OPF method impractical.
Figure 6: The visualization effectively emphasizes the active power of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method.
Figure 6: The visualization effectively emphasizes the active power of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method.
Figure 7: The visualization effectively emphasizes the economic cost of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method.
4.5 Visualization Analysis
To better understand the superiority of the proposed Powerformer, we present illustrations that depict active power and economic cost across the 10 transmission sections.
Active Power Comparison. By comparing the active power of all scenario samples before and after Powerformer adjustment, across all ten transmission sections on the IEEE 118-bus system, we can find that all test samples are successfully adjusted into the pre-determined range. Among these cases shown in Fig. 7, the majority involve adjusting the situation where the power flow exceeds the limit to align with the range of the upper bound, thereby guaranteeing the safe operation of the power system and ensuring its long-term stability.
Economic Cost Comparison. By comparing the economic cost of all scenario samples before and after Powerformer adjustment on all ten transmission sections of the IEEE 118-bus system, we can find that the economic cost of all test samples has been reduced. As shown in Fig. 7, after undergoing further adjustment through Powerformer, the economic cost of nearly every sample has been effectively reduced, with no instances of cost increase. Hence, it can be inferred that our method reliably aims to reduce economic cost.
Figure 8: The performance of our proposed Powerformer and its ablations on the IEEE 118-bus system.
4.6 Ablation Study
By conducting thorough ablation studies and performing a detailed analysis, we go deeper insights into our proposed Powerformer (PF) architecture. We carefully assess and quantify the contributions of each component, revealing how they individually impact the overall effectiveness and superiority of the algorithm. Fig. 8 presents the comparison results obtained from an array of ablation experiments conducted on IEEE 118-bus system with 10-section task. These research findings provide valuable insights, clarifying the roles of each component in improving performance.
Powerformer-E (PF-E). This method only employs GNN to learn state representation, without performing state factorization operation and integrating transmission section information. This method lacks specific feature extraction for individual electrical factors, resulting in the coupling of features. Moreover, the absence of section information undermines the accuracy of the representation vector in depicting the current system state, thereby impeding effective decision-making.
Powerformer-S (PF-S). This method only utilizes the representation of Value component $\boldsymbol{V}$ (Fig. 2) as the final representation of the system state. In contrast to Powerformer-E, it involves the addition of an additional factorization operation. By analyzing the results in Fig. 8, it is evident that comparing Powerformer-S with Powerformer highlights the inadequacy of relying solely on state representations. This emphasizes the importance of incorporating section information.
Powerformer-M (PF-M). This method skips the factorization operation and use the raw four-dimensional electrical features as input to learn the final representation within the Powerformer architecture. In contrast to Powerformer-E, it incorporates an additional operation for combining the transmission section information. The less satisfactory result in Fig. 8 implies that the disentangling operation plays a crucial role in determining the final performance.
5 Conclusion
In this paper, we propose and present a specialized Transformer architecture, named Powerformer, which integrates the transmission section power flow information in order to address the issue of power flow adjustment. The utilization of our proposed architecture facilitates controlled power dispatch, leading to improved stability of the power system. Specifically, our proposed Powerformer incorporates transmission section information into the existing Transformer and develops a section-adaptive attention mechanism. Furthermore, our approach disentangles power system graph data into independent electrical factor graphs to conduct a multi-factor attention mechanism, and employs graph neural networks to extract system state features while considering the graph topological structure of the power system. This allows for the robust representations of the sectional power system state, enabling the development of efficient strategies to effectively manage and adjust transmission section power flow. In our future work, we will focus on developing more efficient architectures for factorization and attribution to enhance the performance in adjusting transmission section power flow.